### v1

# Observe Then Act: Asynchronous Active Vision-Action Model for Robotic Manipulation

Abstract
In real-world scenarios, many robotic manipulation tasks are hindered by occlusions and limited fields of view, posing significant challenges for passive observation-based models that rely on fixed or wrist-mounted cameras. In this paper, we investigate the problem of robotic manipulation under limited visual observation and propose a task-driven asynchronous active vision-action model.Our model serially connects a camera Next-Best-View (NBV) policy with a gripper Next-Best Pose (NBP) policy, and trains them in a sensor-motor coordination framework using few-shot reinforcement learning. This approach allows the agent to adjust a third-person camera to actively observe the environment based on the task goal, and subsequently infer the appropriate manipulation actions. We trained and evaluated our model on 8 viewpoint-constrained tasks in RLBench. The results demonstrate that our model consistently outperforms baseline algorithms, showcasing its effectiveness in handling visual constraints in manipulation tasks. Our code and demonstration videos are available at https://hymwgk.github.io/ota/.
Index Terms:
Reinforcement learning , Robotic manipulation, Active vision, Next-best-view, Sensorimoter.
I Introduction
Despite the reinforcement learning based models for robotic manipulation can learn to directly infer effective actions from camera-provided observation images through interaction with the environment, exhibiting strong scene perception and reasoning capabilities, they still heavily rely on cameras positioned based on prior knowledge, such as third-person fixed cameras and first-person wrist-mounted camera[1][2][3]. While these fixed third-person cameras and wrist-mounted cameras help reduce the sample variance of observations during interaction[3][4], they often struggle with providing sufficient visual information when dealing with manipulation scenarios under occlusions and limited observations due to unadjustable viewpoint positions or fixed camera-gripper configuration.
Humans can independently adjust their head position and eye fixation based on their understanding of task goals and the current environment state to obtain the most advantageous visual information for making interactive action decisions[5]. Inspired by this, recent works in active vision-action joint learning[3][6][7][8] aim to mimic human capabilities by enabling agents to actively adjust their sensory “eye” viewpoints and interact with the world using their motor “hands.”
Similar to humans, the active vision-action model exists in two categories: synchronous and asynchronous[10]. The synchronous vision-action model makes real-time decisions on the next visual sensor and motor actions based on current observation information[7][8][11][12]. This setup requires the model to effectively coordinate sensor motor actions in the next moment and to infer motor actions even when sufficient observation information may not be available which impose higher demands on the model’s dynamic modeling capabilities, usually requiring more interaction steps and making it difficult to apply to manipulation tasks requiring large viewpoint changes.
Although humans are capable of making synchronous decisions, hand movements and changes in viewpoint typically do not occur simultaneously; instead, we generally select the optimal eye viewpoint before executing an interaction. Similarly, the asynchronous vision-action model separates visual sensing and motor actions over time, allowing the vision policy to focus less on coordinating both actions simultaneously when determining the observation viewpoint. Instead, the model can focus more on task-specific manipulations and distribute the sensor-motor coordination throughout the entire task cycle. Although such models are currently more often applied to object grasping tasks in occluded scenes[13][14], we believe that the asynchronous approach may be more suitable for common sequential manipulation tasks.
In this paper, we further explore the problem of robotic manipulation under limited observation and introduce a task-driven asynchronous active vision-action model. The model is designed as a sequential dual-agent structure, as illustrated in Fig.LABEL:fig:head_image. It consists of two agents: a Next-Best-View (NBV) agent for inferring optimal camera viewpoints, and a Next-Best-Pose (NBP) agent for determining the gripper’s action based on the observations from the NBV-inferred viewpoint. By alternating between sensor and motor action inference in each episode, our model enables active perception of the environment and effective interaction with the environment to achieve the manipulation task objectives.
We train our model using a few-shot reinforcement learning approach[2] and introduce a viewpoint-aware demo augmentation method to maximize the utilization of limited demonstration trajectories. To address the challenge posed by the expanded observation-action space resulting from the active viewpoint, we employ a viewpoint-centric voxel alignment technique, which rotates the scene voxel to align with the camera viewpoint. To further stabilize the training process, we incorporate a combined task-agnostic auxiliary reward which provides sufficiently task-aligned learning signals throughout the training process, encouraging the agent to interact effectively with the environment and select viewpoints that offer richer visual information for more informed gripper action decisions. We conduct comparison experiments and analyses on 8 manipulation tasks with limited effective viewpoints in RLbench[9]. The results demonstrate our model’s effectiveness in selecting task-relevant viewpoints and its robustness in inferring manipulation actions based on the current state and task requirements. The main contributions of this paper are as follows:
  • We propose a task-driven asynchronous active vision-action model that uses a novel dual-agent system to achieve effective robotic manipulation under limited observation.
  • We introduce methodological innovations, including viewpoint-aware demo augmentation, viewpoint-centric voxel alignment, and a task-agnostic auxiliary reward strategy, to enhance model performance, learning efficiency, and training stability.
  • Comprehensive experiments on public RLbench tasks validate the effectiveness and robustness of our approach.
II Related Work
Active Vision for Manipulation. Active vision aims to enable agents to actively observe the environment based on specific task goals. It has been widely applied in tasks such as indoor visual navigation[12][15], object searching[16, 17, 18, 19] and 3D reconstruction[20]. In robotic manipulation, active vision is primarily used for grasping pose detection in occluded environments[7][13][14][21][22] and for detecting fruit nodes in automated harvesting[23, 24, 25]. Recent studies have extended the use of active vision to more general robotic manipulation tasks. For instance, Cheng et al[11] proposed an active vision model with a target detector module to handle object-pushing task in occluded scenes. Lv et al[8] introduced a model-based sensing-aware framework for manipulating flexible objects. Matthew et al[4] employed a self-prediction approach to learn viewpoint adjustments while simultaneously developing motion skills for robotic arm joints. Similarly, Shang et al[6] utilized a self-supervised reward mechanism to jointly learn visual sensing and motor skills. Unlike these methods, our active vision-aciton model does not require a task-specific target detector and employs an RL-based dual agent asynchronous decision-making structure.
3D Voxel-based Manipulation. 3D voxels are highly effective in representing spatial structures[26], making them widely used in domains such as medical imaging[27][28] and autonomous driving[29][30]. Recently, they have also been adopted in robotic manipulation models for visual representation. For instance, James et al.[2] transformed original RGB-D images into 11-dimensional voxel grids and developed a 3DUnet-based reinforcement learning model to extract scene features and infer manipulation actions. Similarly, Shridhar et al.[31] used discrete 3D voxel images as input, employing the Perceiver network[32] as a backbone for end-to-end behavioral cloning, trained on demonstration trajectories. Leveraging the strong spatial structure representation and robust structural invariance of voxels under viewpoint changes—which aids in viewpoint learning—our model also uses voxelized scene images as input.
Figure 2: Comparison of regular RL and asynchronous active vision-action decision processes. We introduce an intermediate time step $t^{\prime}$ between $t$ and $t+1$, dividing the interaction into NBV policy for viewpoint selection and the NBP policy for action selection. The task reward $r_{\text{t'}}$ for state $\mathbf{s}$ is derived from the reward $r_{\text{t+1}}$ after the NBP interaction, which enables sensor-motor joint training through shared task rewards.
Attention-driven Robotic Manipulation. The attention mechanism, widely applied in fields such as object detection and natural language processing, contributes to robotic manipulation by focusing on the most task-relevant scene information from raw image inputs, improving both training efficiency and model performance[33]. ARM [1] and C2FARM [2] use attention mechanisms to focus on key visual input regions, with ARM selecting noteworthy 2D regions via DQN, while C2FARM predicts regions of interest (ROIs) in 3D space. Liang et al. [34] pre-train a scene saliency attention encoder offline using human-annotated saliency maps and then train the reinforcement learning operation agent online based on the scene features output by this encoder. However, these methods largely rely on single or multiple fixed cameras from different viewpoints to capture images of the scene, lacking the capability to actively choose their viewpoints. Our model partially incorporates the attention mechanism from C2FARM and endows the model with the ability to predict the optimal task-related observation viewpoint, further mimicking the human’s active visual perception capability.
III background
Figure 3: Overview of our model pipeline. The NBV policy infers the scene ROI position $\mathbf{f}^{*}_{t}$ based on the current observation and determines the optmal viewpoint $\mathbf{v}^{*}$ for observing the ROI to achieve the task goal. The NBP agent infers the gripper actions based on the scene ROI image from viewpoint $\mathbf{v}^{*}$ and the proprioceptive state.
Problem Formulation. We divide one asynchronous vision-action time step $t\sim t+1$ into the viewpoint decision and transform process of the NBV policy from $t\sim t^{\prime}$ and the gripper action decision and interaction process of the NBP policy from $t^{\prime}\sim t+1$. Therefore the asynchronous active vision manipulation problem can be formalized as a POMDP described by the tuple $<\mathcal{O},\mathcal{A^{\text{c}}},P,\mathcal{A^{\text{g}}},\mathcal{O^{\prime}},P^{\prime},R,\gamma>$ As shown in Fig.2, $\mathcal{O}$ and $\mathcal{O^{\prime}}$ represent the observation spaces of the agent at times $t$ and $t^{\prime}$, respectively. The NBV policy $\pi_{\text{v}}$ determines the camera viewpoint action $\mathbf{a}_{t}^{\text{c}}=\pi_{\text{v}}(\cdot|\mathbf{o}_{t})\in\mathcal{A^{\text{c}}}$ based on the observation $\mathbf{o}_{t}\in\mathcal{O}$, and obtains a new observation $\mathbf{o}_{t^{\prime}}\in\mathcal{O^{\prime}}$ through transition probability $p(\mathbf{o}_{t^{\prime}}|\mathbf{o}_{t},\mathbf{a}_{t}^{\text{c}})\in P$. The NBP policy $\pi_{\text{g}}$ then determines the gripper action $\mathbf{a}_{t^{\prime}}^{\text{g}}=\pi_{\text{p}}(\cdot|\mathbf{o}_{t^{\prime}})\in\mathcal{A^{\text{g}}}$ based on the new observation $\mathbf{o}_{t^{\prime}}$. Through the transition probability $p^{\prime}(\mathbf{o}_{t+1}|\mathbf{o}_{t^{\prime}},\mathbf{a}_{t^{\prime}}^{\text{g}})\in P^{\prime}$, scene observation $\mathbf{o}_{t+1}$ can be obtained by camera at time $t+1$, and the environment provides the task reward $r_{t+1}\in R$. Our goal is to jointly learn the policies $\pi^{*}_{\text{v}}$ and $\pi^{*}_{\text{g}}$ that maximize the expected sensor-motor joint reward for the given manipulation task, which is $\pi^{*}_{\text{v}},\pi^{*}_{\text{g}}=\arg\underset{\pi_{\text{v}},\pi_{\text{p}}}{\max}\mathbb{E}\Big{[}\sum_{t=0}^{\infty}\gamma^{t}R(\mathbf{o}_{t},\mathbf{a}^{\text{c}}_{t},\mathbf{a}^{\text{g}}_{t^{\prime}})\Big{]}$.
Observation and Action Space We define the observation spaces for the NBV and NBP policies as identical, with $\mathcal{O}=\mathcal{O^{\prime}}=\{\mathcal{O}^{\text{v}},\mathcal{O}^{\text{p}}\}$, where $\mathbf{o}^{\text{v}}\in\mathcal{O}^{\text{v}}$ represents the 3D voxel grid voxelized from the scene RGB-D image captured by the camera, and $\mathbf{o}^{\text{p}}\in\mathcal{O}^{\text{p}}$ is the proprioceptive state. Specifically, the raw RGB-D image $\mathbf{o}^{\text{i}}$ is converted into a point cloud $\mathbf{o}^{\text{pc}}\in\mathbb{R}^{n\times 6}$ using the function $\mathbf{P}(\mathbf{o}^{\text{i}}):\mathbf{o}^{\text{i}}\rightarrow\mathbf{o}^{\text{pc}}$, which is then voxelized through function $\mathbf{V}(\mathbf{o}^{\text{pc}},\mathbf{e},\mathbf{c}):\mathbf{o}^{\text{pc}}\rightarrow\mathbf{o}^{\text{v}}$, where $\mathbf{e}$ denotes the voxel resolution and $\mathbf{c}$ represents the voxel center coordinates. The resulting 3D voxel grid $\mathbf{o}^{\text{v}}\in\mathbb{R}^{n\times(3+M+1)}$ contains the center coordinates, $M$-dimensional features, and a occupancy flag for each voxel. The proprioceptive state $\mathbf{o}^{\text{p}}=\{\mathbf{v},\mathbf{g}\}$ comprises the camera viewpoint position $\mathbf{v}\in\mathbb{R}^{3}$ and the gripper pose $\mathbf{g}\in\mathbb{R}^{6}$.
Following common settings[14][21][22], we restrict the camera viewpoint’s movement to the upper hemisphere at a fixed radius of $r^{\text{v}}=1.2\text{m}$ centered on the table, illustrated in Fig.8. The camera’s optical axis is assumed to always be directed towards the table center thus the viewpoint position can be described by spherical coordinates $\mathbf{v}=(r^{\text{v}},\theta^{\text{v}},\phi^{\text{v}})$. We than define the NBV policy’s action as $\mathbf{a}^{\text{c}}=\{\mathbf{k}^{\text{v}},\mathbf{k}^{\text{f}}\}\in\mathcal{A}^{\text{c}}$, where $\mathbf{k}^{\text{v}}\in\mathbb{R}^{2}$ represents the discrete viewpoint index for $(\theta^{\text{v}},\phi^{\text{v}})$, and $\mathbf{k}^{\text{f}}\in\mathbb{R}^{3}$ denotes the discrete center position index of a region of interest(ROI) that requires attention within the scene. The NBP policy’s action is defined as $\mathbf{a}^{\text{g}}=\{\mathbf{k}^{\text{g}},d\}\in\mathcal{A}^{\text{g}}$, where $\mathbf{k}^{\text{g}}$ represents the discrete 6-DoF gripper pose index, and $d$ is the goal distance between the parallel gripper tips.
IV Method
In this section, we provide a detailed explanation of the fundamental structure of our active vision-action model, along with three innovative strategies designed to enhance both model performance and learning efficiency.
IV-A Model Structure
Fig3 illustrates the basic pipline of our model. At time $t$, the voxelized image observation and proprioceptive state undergo a coordinate transformation for viewpoint alignment (Section IV-B), resulting in the processed observation $\mathbf{o}_{t}$. The NBV policy network, parameterized by $\Theta$, uses $\mathbf{o}_{t}$ to determine the discrete position index $\mathbf{k}^{\text{f}}$ of the ROI center in the scene, as described in Eq.(1). Simultaneously, the network infers the corresponding discrete index $\mathbf{k}^{\text{v}}$, representing the viewpoint that would best improve task success. The camera is then adjusted to the goal viewpoint $\mathbf{v}^{*}$, which is derived from the selected $\mathbf{k}^{\text{v}}$.
$\mathbf{a}_{t}^{\text{c}}=\arg\underset{\mathbf{k}^{\text{v}},\mathbf{k}^{\text{f}}}{\max}Q_{\Theta}(\mathbf{o}_{t},\mathbf{k}^{\text{v}},\mathbf{k}^{\text{f}})$ (1)
Based on the predicted ROI position $\textbf{f}^{*}$, derived from $\mathbf{k}^{\text{f}}$, the raw image observation from the viewpoint is cropped and aligned to generate the corresponding ROI 3D voxel grid $\mathbf{o}^{\text{v}}_{t^{\prime}}$. This enables us to construct the observation $\mathbf{o}_{t^{\prime}}=\{\mathbf{o}^{\text{v}}_{t^{\prime}},\mathbf{o}^{\text{p}}_{t^{\prime}}\}$. Using the NBP policy network parameterized by $\Phi$, we infer the gripper closure action $d$ and the discrete gripper pose index $\mathbf{k}^{\text{g}}$ which is then transformed into the goal pose $\mathbf{g}^{*}$. Finally, a complete sensor-motor interaction step is achieved after gripper exectuation and rewarded by $r_{t+1}$.
$\mathbf{a}_{t^{\prime}}^{\text{g}}=\arg\underset{\mathbf{k}^{\text{g}},d}{\max}Q_{\Phi}(\mathbf{o}_{t^{\prime}},\mathbf{k}^{\text{g}},d)$ (2)
We denote the sub-transition for the NBV policy as $\text{T}_{t}^{\text{v}}=<\mathbf{o}_{t},\mathbf{a}_{t}^{\text{c}},\mathbf{o}_{t^{\prime}}>$ and for the NBP policy as $\text{T}_{t}^{\text{p}}=<\mathbf{o}_{t^{\prime}},\mathbf{a}_{t^{\prime}}^{\text{g}},\mathbf{o}_{t+1}>$. Since the NBV and NBP policies share the same reward $r_{t+1}$, a complete interaction transition is denoted as $\text{T}_{t}=\{\text{T}_{t}^{\text{v}},\text{T}_{t}^{\text{p}},r_{t+1}\}$. Upon completing an interaction episode, the collected transition is stored in the replay buffer together with the demonstration transitions. During model updates, both the NBV and NBP policies are updated using the same reward, thereby achieving sensor-motor joint learning. We employ a 3D-CNN based U-Net[35] as the image encoder for both policy networks, and three-layer fully connected network (FCN) for decoding the viewpoint and gripper actions.
[…]
IV-B Viewpoint-centric Voxel Alignment
Viewpoint movement expands the image observation space, increasing the variance of interaction samples[3]. Specifically, images captured from a static viewpoint are primarily influenced by the scene state, following the mapping $f^{\text{static}}:\mathcal{S}\rightarrow\mathcal{O}^{\text{v}}$. In contrast, active viewpoint movement introduces an additional influence from the viewpoint itself, resulting in the mapping $f^{\text{active}}:\mathcal{S}\times\mathcal{V}\rightarrow\mathcal{O}^{\text{v}}$. Since the 3D-CNN lacks spatial rotation invariance, voxelizing the scene using the world coordinate system as the reference frame can cause features of similar structures from different viewpoints to appear distant in the feature space after encoding, as shown in Fig.4. This misalignment negatively impacts the overall training efficiency of the model.
[…]
$\mathbf{o}^{\text{v}}=\mathbf{V}\big{(}(^{\text{L}}{\text{T}}_{\text{W}}\times\mathbf{o}^{\text{pc}}),\mathbf{e},\mathbf{c}\big{)}$ (3)
[…]
$\mathbf{v}^{*},\mathbf{f}^{*},\mathbf{g}^{*}=^{\text{W}}{\text{T}}_{\text{L}}\times(\mathbf{v}^{*}_{\text{L}},\mathbf{f}^{*}_{\text{L}},\mathbf{g}^{*}_{\text{L}})$ (4)
IV-C Viewpoint-aware Demo Augmentation
Figure 5: Demo trajectory augmentation.For each demo trajectory, we uniformly sample observations (represented by hatched circles) between both viewpoint key frames (solid blue circles) and gripper key frames (solid green circles) to generate additional transition samples, enabling viewpoint-aware demonstration augmentation.
To provide early guidance and reduce the need for extensive exploration, we manually alternate moving the camera and gripper to generate and collect a small number of demonstration trajectories as high-quality examples. Each demo trajectory $\tau_{\text{d}}=\{(\mathbf{o}_{i},\mathbf{a}_{i})\}^{N}$ consists of “observation-action” pairs collected at a fixed sampling rate. Existing demonstration augmentation methods[1][2][31] typically involve two main processes: key frame discovery and observation sampling. In the key frame discovery process, a limited set of key frame indices is selected from a trajectory $\tau_{d}$, and transition samples are constructed based on the corresponding “observation-action” pairs. The observation sampling process then generates additional transition samples by resampling the initial observations between the identified key frames. However, existing methods do not account for viewpoint movement. Our method incorporates key frame discovery and observation resampling, taking into account viewpoint movement, enabling viewpoint-aware demonstration augmentation. Specifically, we define the key frame discovery function as $KF:\tau_{\text{d}}\rightarrow\big{\{}(k_{j}^{\text{c}},k_{j}^{\text{g}})\big{\}}^{M}\subset\mathbb{R}_{+}^{2}$, which retrieves the key frames of both viewpoint and gripper movements in the demonstration trajectory based on predefined rules.
The discovery of gripper key frames $k_{j}^{\text{g}}$ follows the method described in [1]. We define key viewpoints as the first stable viewpoint between every two gripper key frames, corresponding to the frame index $k_{j}^{\text{c}}$. The relationship between key frame indices satisfies $10\leq k_{j-1}^{\text{c}}\leq k_{j-1}^{\text{g}}\leq k_{j}^{\text{c}}\leq k_{j}^{\text{g}}\leq N-$ 1. As shown in Fig.5, after the key frame discovery process, we construct the raw transition samples $\text{T}_{j}$ formalized in Eq.(5).
$\begin{split}\text{T}_{j}&=\{\text{T}_{j}^{\text{v}},\text{T}_{j}^{\text{p}},r_{j}\}\\ &=\{<\mathbf{o}_{k_{j-1}^{\text{g}}},\mathbf{a}_{k_{j-1}^{\text{g}}}^{\text{c}},\mathbf{o}_{k_{j}^{\text{c}}}>,<\mathbf{o}_{k_{j}^{\text{c}}},\mathbf{a}_{k_{j}^{\text{c}}}^{\text{g}},\mathbf{o}_{k_{j}^{\text{g}}}>,r_{j}\}\end{split}$ (5)
To generate additional augmented transitions, we uniformly sample within the intervals $k_{j-1}^{\text{g}}\sim k_{j}^{\text{c}}$ and $k_{j}^{\text{c}}\sim k_{j}^{\text{g}}$, obtaining the augmented frame indices$(k_{j-1}^{\text{g}})^{\text{aug}}$ and $(k_{j}^{\text{c}})^{\text{aug}}$. The resulting augmented transition $\text{T}_{j}^{\text{aug}}$ is formalized in Eq.(6).
$\begin{split}\text{T}_{j}^{\text{aug}}&=\{(\text{T}_{j}^{\text{v}})^{\text{aug}},(\text{T}_{j}^{\text{p}})^{\text{aug}},r_{j}\}\\ &=\{<\mathbf{o}_{(k_{j-1}^{\text{g}})^{\text{aug}}},\mathbf{a}_{(k_{j-1}^{\text{g}})^{\text{aug}}}^{\text{c}},\mathbf{o}_{(k_{j}^{\text{c}})^{\text{aug}}}>,\\ &<\mathbf{o}_{(k_{j}^{\text{c}})^{\text{aug}}},\mathbf{a}^{\text{g}}_{(k_{j}^{\text{c}})^{\text{aug}}},\mathbf{o}_{k_{j}^{\text{g}}}>,r_{j}\}\end{split}$ (6)
Both $\text{T}_{j}$ and $\text{T}_{j}^{\text{aug}}$ are ultimately stored in the replay buffer as early training experiences for the model.
Figure 6: Viewpoint adjustment and ROI selection in a 2D simplified schematic scene. We encourage the model to select non-empty ROIs (green dashed boxes) and maximize the number of occluded voxels within the ROI that can be observed from the adjusted viewpoint $\mathbf{v}_{t^{\prime}}$ compared to the original viewpoint $\mathbf{v}_{t}$. This strategy promotes active interaction with the environment and enables the model to capture more structural information within the ROI.
IV-D Task-agnostic Auxiliary Reward
We use the task completion signal as a sparse task reward to avoid designing targeted rewards for different tasks; however, sparse reward only provides positive feedback to the model at the end of an episode, making it difficult to efficiently guide the model in sensor-motor joint learning. To address this, we introduce two task-agnostic auxiliary rewards that provide real-time feedbacks to the training process.
To encourage effective interaction with the environment and accurate environment modeling, we propose the Interaction Encouragement Reward, denoted as $r_{\text{i}}$ As illustrated in Fig.6, reward $r_{\text{i}}$ provides a positive signal when the NBV policy infers an ROI in a reachable and non-empty area of the scene, and penalizes actions that select an ROI in an empty or unreachable area due to distance or trajectory planning failures. The specific value assignment for $r_{\text{i}}$ is detailed in Eq.(7).
$r_{\text{i}}=\begin{cases}0.02,&\text{ROI reachable and non-empty}\\ 0.0,&\text{ROI reachable but empty}\\ -0.02,&\text{ROI not reachable and empty}\end{cases}$ (7)
To encourage the model to select viewpoints that capture more structural information from the ROI, we introduce the Entropy Reduction Reward , denoted as $r_{\text{e}}$, inspired by [24]. The core idea is to reduce the occupancy entropy within the ROI after adjusting the viewpoint, as illustrated in Fig.6. Specifically, assuming the ROI contains $N$ voxels, the occupancy entropy of the $i$-th voxel is calculated using the Shannon entropy function, as shown in Eq.(8), where $p_{\text{o}}(i)$ represents the voxel’s occupancy probability. If a voxel is occluded, its occupancy probability is set to $p_{\text{o}}=0.5$.
$e_{i}=-p_{\text{o}}(i)\log_{2}\big{(}p_{\text{o}}(i)\big{)}-\big{(}1-p_{\text{o}}(i)\big{)}\log_{2}\big{(}1-p_{\text{o}}(i)\big{)}$ (8)
The total occupancy entropy change of the ROI from time $t$ to $t^{\prime}$ is expressed as:
$\Delta E=\frac{1}{N}\sum_{i=1}^{N}e^{t}_{i}-\frac{1}{N}\sum_{i=1}^{N}e^{t^{\prime}}_{i}=E_{t}-E_{t^{\prime}}$ (9)
We define the reward $r_{\text{e}}$ as Eq.(10), where the coefficient $k$ is empirically set to 1.0.
$r_{\text{e}}=k\cdot\Delta E,k=1.0$ (10)
The auxiliary reward $r_{\text{i}}$ necessitates the cooperation of both the NBV and NBP policies, whereas $r_{\text{e}}$ is exclusive to the NBV policy. Therefore, the reward for updating the NBP policy is expressed as $r_{\text{nbp}}=r_{\text{task}}+r_{\text{i}}$, while the reward for updating the NBV agent is $r_{\text{nbv}}=r_{\text{task}}+r_{\text{e}}+r_{\text{i}}$, where $r_{\text{task}}\in[0,1]$ is determined by task completion. It is important to note that the introduced rewards are designed to be task-agnostic, which ensures that the rewards can generalize across different tasks without the need for manual tuning.
Figure 7: Reletive task occlusion rate of selected tasks. We use the active viewpoint occlusion rate of 30 demonstration trajectories for each task as a benchmark. Based on this, we calculate the relative task occlusion rates for the front, overhead, and wrist viewpoints in the demonstration trajectories of each task.
V Experiment Setup
V-A Task Setting
We trained and evaluated our model on RLBench [9], a public benchmark designed for reinforcement learning manipulation algorithms. When selecting evaluation tasks, we considered two primary criteria: 1) Single task: each task must have at least one fixed viewpoint with a high level of visual occlusion. 2) Task range: no single fixed or wrist-mounted camera can provide a low occlusion rate for all tasks. Since the region of interest (ROI) the robot focuses on may change during different manipulation phases, and there may be no clear manipulation objects (e.g., when the gripper is transferring in midair), it is challenging to use an image segmentor to locate objects and define the task occlusion rate as in previous methods[14][24]. To address this, we define the Task Occlusion Rate (TOR) for a given viewpoint as the average occupancy entropy of the ROI where the gripper’s goal actions occur, based on observations from that viewpoint across $L=100$ demo trajectories, as described in Eq.(11), Here, $E^{l}_{t^{\prime}_{k}}$ represents the ROI occupancy entropy at the $k$-th keyframe in trajectory $l$ after viewpoint adjustment where the ROI center is set to the gripper’s goal position in the demo observation. To compare the occlusion levels across different tasks, we assume that the collected demo trajectories contain the strongest task-specific prior knowledge. Thus, we can obtain the Relative TOR by dividing the TOR of the selected viewpoint by the TOR of the human-demoed active viewpoint (oracle).
$\text{TOR}=\frac{1}{L}\sum_{l=1}^{L}\sum_{k=0}^{K_{l}}E^{l}_{t_{k}^{\prime}}$ (11)
Fig.7 presents the relative TOR statistics for 8 selected RLbench tasks under common overhead, front , wrist-mounted cameras as well as the human-demoed active camera, based on $100$ demo trajectories. The results show that, due to task characteristics and initial scene randomization, both fixed and wrist viewpoints experience varying degrees of occlusion across the 8 tasks. In contrast, the active camera demonstrates the potential to achieve lower task occlusion.
V-B Baselines
We compaired our model against two categories baseline models: 1) visual manipulaiton model that utilze common used static third-person cameras and first-person wrist-mounted camera; 2) Active third-person camera based manipulation models.
For the first category, we compared our active vision-action model against the recent visual-based C2FARM [2] model with static front, overhead camera as well as the wrist-mounted camera settings. This comparison aims to explore the potential effectiveness of our active vision-action model in handling tasks with limited observability.
For the second category, using the same third-person active camera setting, we firstly compared our model against the Behavior Cloning (BC) model. Through this comparison, we seek to investigate whether our model not only mimics the given limited demonstrations but also learns new sensor-motor skills through interaction with the environment. Additionally, we tested a Random viewpoint model, following [3], which shares the same architecture as ours but with the learning of the NBV policy network disabled (learning rate set to 0). This setup allows us to evaluate the importance of the active viewpoint selection in the NBV policy and to assess how much performance improvement is achieved through active viewpoint adjustment as opposed to random viewpoint choices.
V-C Training Setting
We first collected 30 demo trajectories for each task and stored them in the replay buffer as prior samples or the training dataset, and all models were trained from scratch.
The implementation of C2FARM was based on their official release and trained using sparse task rewards. The Behavior Cloning model was trained with the Mean Squared Error (MSE) loss, while the Random model utilized the same reward structure as our model, as described in SectionIV-D. All models were trained using the Adam optimizer with an initial learning rate of $0.0005$, with no learning rate decay applied during training. Each model was trained for 40,000 updates with a batch size of 128, and all experiments were conducted on a machine equipped with 4 NVIDIA GTX 4090 GPUs.
TABLE I: Task success rate: Our model outperform baselines on most tasks and on average . Mean and std of 100 tests are reported.
Models Basketball in Hoop Meat off Grill Meat on Grill Open Drawer Press Switch put rubbish in bin Take Money Out Safe Unplug Charger On Average
C2FARM(wrist) 0.93±0.02 0.01±0.01 0.11±0.02 0.44±0.07 0.90±0.02 0.00±0.00 0.52±0.01 0.30±0.02 0.40±0.35
C2FARM(overhead) 0.87±0.02 0.00±0.00 0.28±0.05 0.03±0.01 0.89±0.03 0.61±0.04 0.24±0.03 0.89±0.04 0.48±0.36
C2FARM(front) 0.26±0.03 0.28±0.04 0.68±0.04 0.37±0.04 0.86±0.06 0.91±0.02 0.66±0.06 0.00±0.00 0.50±0.30
Random(active) 0.00±0.00 0.00±0.00 0.00±0.00 0.00±0.00 0.00±0.00 0.29±0.02 0.00±0.00 0.00±0.00 0.04±0.09
BC(active) 0.00±0.00 0.00±0.00 0.02±0.01 0.19±0.03 0.51±0.02 0.00±0.00 0.38±0.03 0.02±0.01 0.14±0.19
Ours(active) 0.81±0.02 0.56±0.04 0.73±0.04 0.50±0.03 0.85±0.03 0.96±0.01 0.78±0.02 0.92±0.01 0.76±0.15
TABLE II: Task episode length. The maximum length of each task episode is limited to 10 steps. Mean and std of 100 tests are reported.
Models Basketball in Hoop Meat off Grill Meat on Grill Open Drawer Press Switch put rubbish in bin Take Money Out Safe Unplug Charger On Average
C2FARM(wrist) 3.73±0.18 9.96±0.04 9.27±0.20 7.06±0.49 3.64±0.23 9.98±0.03 6.39±0.11 7.75±0.17 7.22±2.39
C2FARM(overhead) 5.21±0.17 10.00±0.01 8.50±0.28 9.78±0.07 4.13±0.23 6.52±0.21 8.46±0.21 3.93±0.22 7.07±2.30
C2FARM(front) 8.43±0.15 8.66±0.18 5.91±0.24 8.04±0.20 4.49±0.40 3.22±0.13 5.63±0.41 9.99±0.02 6.80±2.20
Random(active) 10.00±0.00 10.00±0.00 10.00±0.00 10.00±0.00 10.00±0.00 7.76±0.27 10.00±0.00 10.00±0.00 9.72±0.75
BC(active) 10.00±0.00 10.00±0.00 9.93±0.05 8.68±0.24 6.74±0.28 10.00±0.00 7.64±0.27 9.90±0.07 9.11±1.22
Ours(active) 5.35±0.08 6.43±0.28 5.15±0.26 6.65±0.19 3.89±0.28 3.33±0.09 4.30±0.17 3.91±0.07 4.88±1.16
VI Result
Main Performance. We first analyzed the average task success rate and the average episode length across different algorithms, which provide a macro-level assessment of the models’ proficiency in learning sensor-motor coordination.
Table I presents the average success rate for tasks across 100 testing episodes under 3 evaluation seeds. Our model achieved the highest average success rate across the eight tasks, outperforming the second-best by 21%. It also exhibited relatively high manipulation performance in scenarios with significant occlusion from fixed viewpoints, such as meat_on_grill, meat_off_grill and unplug_charger.
Similarly, Table II shows that the average episode length for tasks follows the same trend as the success rate. Our model achieved the shortest average manipulation steps across all tasks, outperforming the second-best model by 28.2%. Furthermore, it demonstrated the lowest variance in manipulation steps, highlighting the model’s high proficiency and consistency after training.
Moreover, we observed that in tasks where the occlusion rate is low or the scene object positions are relatively predictable, fixed and wrist-mounted cameras benefit from smaller variance in observation samples, leading to enhanced skill acquisition. For example, in tasks such as basketball_in_hoop, press_switch, put_rubbish_in_bin, the C2FARM model achieved comparable success rates from both front and overhead views, occasionally even surpassing our model. However, it is important to note that when these fixed cameras fail to capture adequate observations, performance tends to decline sharply, leading to higher performance variance across the 8 tasks.
In contrast, our model is less sensitive to visual occlusion and consistently maintains strong manipulation performance. Under the active viewpoint setting, although the BC model demonstrated some ability to adjust its viewpoint (as detailed below), its effectiveness was constrained by the limited number of training samples, which restricted its capacity to learn interaction skills. As a result, it frequently failed to complete the assigned tasks, achieving an average success rate of only 14%. Meanwhile, the random viewpoint model performed poorly across all 8 tasks, exhibiting the weakest manipulation performance, indicating its inability to effectively learn sensor-motor coordination skills.
Figure 8: Viewpoint distribution overview: Red circles indicate the distribution of manually set viewpoints across 30 demo trajectories, while blue triangles represent the distribution of viewpoint positions during successful episodes of our model over 100 test episodes. The model’s inferred viewpoints do not strictly mimic the demonstration trajectories but instead exhibit a degree of refinement and generalization.
Figure 9: Partial rollout cases of our model. The figure presents the image observations corresponding to the first three inferred viewpoints $\mathbf{v}^{*}$ and the associated gripper actions $\mathbf{g}^{*}$ for four representative tasks during the test episodes.
TABLE III: A comparison of information gain to baselines.
      Models              ASSIG       (avg/max)              AIIG       (avg/max)
      C2FARM(wrist)       0.0/0.0       N/A
      C2FARM(overhead)       0.0/0.0       0.0/0.0
      C2FARM(front)       0.0/0.0       0.0/0.0
      Random(active)       0.0/5.1%       0.0/5.2%
      BC(active)       0.7%/2.3%       5.6%/18.1%
      Ours(active)       3.5%/6.9%       6.8%/25.1%
Active Perception Strategy Analysis. We further analyzed the perception strategies of all models during the operation process using two metrics: Average Single Step Information Gain (ASSIG) and Average Initial Information Gain (AIIG). These metrics measure the models’ ability to acquire observation information through multi-step and single-step interactions, respectively. Specifically, we define ASSIG as the average reduction in entropy within the ROI before and after a single viewpoint movement.Similarly, the initial information gain is defined as the average reduction in entropy of the ROI after a viewpoint movement, compared to the initial viewpoint. These two metrics are given by:
$\text{ASSIG}=\frac{1}{L}\sum_{l=1}^{L}\sum_{k=0}^{K_{l}}(E^{l}_{t_{k}}-E^{l}_{t^{\prime}_{k}})$ (12)
$\text{AIIG}=\frac{1}{L}\sum_{l=1}^{L}\sum_{k=0}^{K_{l}}([E^{l}_{t_{k}}]_{\mathbf{v}_{t_{0}}}-E^{l}_{t^{\prime}_{k}})$ (13)
Where $L=100$ denotes the number of testing episodes, $K_{l}$ represents the number of steps required to complete episode $l$, and $E^{l}_{t_{k}}$ is defined in Eq.(9). The term $[E^{l}_{t_{k}}]_{\mathbf{v}{t_{0}}}$ represents the visual entropy of the current ROI from the episode’s initial viewpoint. The statistical results in Table III show that our model significantly improves the observation of the ROI region by independently adjusting camera viewpoints, outperforming the second-best BC model by 2.8% and 1.2% in the ASSIG and AIIG scores, respectively. In contrast, static front and overhead cameras, which are unable to move, fail to achieve either single-step or initial information gain through viewpoint shifts. Similarly, the wrist-mounted camera, being unable to move independently of the gripper, is limited to inferring the next gripper actions based on the current viewpoint. This results in a zero ASSIG score and makes it unsuitable for AIIG computation.
Fig.8 illustrates the distribution of viewpoints learned by our model alongside the key viewpoints distribution from 30 demo trajectories. Notably, the viewpoints learned by our model do not strictly follow the patterns observed in the demonstration trajectories. For tasks such as put_rubbish_in_bin, press_switch, unplug_charger, and meat_off_grill, the learned viewpoints are more concentrated. In contrast, for tasks like meat_on_grill, open_drawer, and take_money_out_safe, the model explores a broader range of viewpoints. This variation indicates that our model has effectively learned to select task-relevant viewpoints through interactive learning, rather than merely replicating the active perception patterns from the demonstration data. Furthermore, Fig. 8 presents camera observations at keyframe moments of partial task episodes, showcasing our model’s capacity for active visual perception during task execution.
Figure 10: Comparison of success, timeout, and execution failure rate across models.
Environmental Modeling and Interactive Capabilities. To further evaluate the environment modeling and interaction capabilities of various models, we recorded the success rate, timeout rate, and execution failure rate over 100 test episodes for each model, as illustrated in Fig.10. The C2FARM model showed a high likelihood (up to 42%) of episodes terminating prematurely due to failures in low-level trajectory execution. Similarly, the random viewpoint model saw the majority of its test episodes end in failure because of execution errors. While execution failures accounted for only 4% of the BC model’s episodes, as many as 82% of the episodes failed due to timeouts. However, our model showed only a 13% rate of episodes ending prematurely due to trajectory execution failures and a mere 9% failure rate due to timeouts. These results suggest that our model effectively integrates low-level motion execution considerations when inferring manipulation actions, resulting in more robust performance across the test episodes.
Additionally, as shown in Table IV, we analyzed the average steps to first interaction with scene objects and the average number of non-interactive steps during testing for each model. Our findings indicate that the behavior cloning exhibited a relatively high number of non-interactive steps, which aligns with the high timeout rate shown in Fig.10. This pattern suggests that behavior cloning struggles to learn precise interactive manipulations from a limited number of demonstration trajectories, leading to frequent non-interactive steps and episode timeouts. In contrast, our model consistently initiates interaction with the environment more quickly and significantly reduces the time spent on non-interactive actions. These results demonstrate that, after training, our model achieves superior accuracy in modeling environmental dynamics and understanding structural relationships.
TABLE IV: First interaction step and non-interaction step comparison.
Models First Interaction Step Non-interaction Step
C2FARM(wrist) 3.21±1.82 4.86±2.70
C2FARM(overhead) 2.97±1.70 4.41±2.77
C2FARM(front) 3.41±2.28 4.17±2.74
Random(active) 9.07±1.47 9.49±1.07
BC(active) 6.31±2.46 7.54±2.15
Ours(active) 1.64±0.61 2.15±1.03
Ablation Study In this section, we investigate the impact of three key strategies on manipulation performance and various evaluation metrics: (1) viewpoint-centric voxel alignment (Section IV-B), (2) viewpoint-aware demon augmentation (Section IV-C), and (3) task-agnostic auxiliary rewards (Section IV-D). To assess the individual and combined contributions of these strategies, we conducted separate ablation tests for each, followed by a comprehensive test removing all three strategies.
As shown in Table V, each of the three strategies contributed to notable improvements in the success rate of our proposed model. Among the three, voxel alignment and viewpoint-aware demonstration augmentation had the most substantial impact, with the success rate dropping by approximately 40% when either of these strategies was ablated, highlighting their critical roles. The auxiliary reward strategy, while less significant than the other two, still provided approximately a 6% improvement, demonstrating its contribution to overall performance.
Beyond success rate improvements, all three strategies contributed to reductions in episode length, overtime ratio, planning failure rate, and the average number of non-interactive steps per episode. This demonstrates their broader importance to the algorithm’s effectiveness. Interestingly, while the voxel alignment strategy significantly boosts overall performance, it leads to a reduction in ASSIG score. We hypothesize that without viewpoint alignment, the NBV agent avoids potential issues with out-of-bounds actions and clipping, which may simplify the learning of viewpoint transformations. However, the overall performance still confirms that viewpoint alignment plays a crucial role in improving operational effectiveness.
When all three strategies were ablated, the performance across the eight tasks showed the most severe decline, with the average success rate dropping to just 22%, even lower than the static-viewpoint baseline. This highlights the synergistic benefits of these three strategies, as their combination significantly boosts the model’s overall performance.
TABLE V: Ablation over voxel alignment,demo augmentation and task-agnostic auxiliary rewards
Models Success Rate Episode Length Overtime Rate Planning Failed Rate
Ours(w/o align) 0.34±0.41 7.61±2.94 0.14±0.12 0.53±0.39
Ours(w/o aug) 0.31±0.37 7.99±2.44 0.15±0.17 0.55±0.37
Ours(w/o aux) 0.70±0.19 5.52±1.51 0.15±0.10 0.16±0.12
Ours (w/o align+aug+aux) 0.22±0.34 8.46±2.41 0.22±0.22 0.57±0.34
Ours 0.76±0.15 4.88±1.16 0.10±0.06 0.15±0.11
Models First Interaction Non-interaction AIIG ASSIG
Ours(w/o align) 4.33±2.80 5.81±3.30 0.046±0.05 0.082±0.07
Ours(w/o aug) 4.49±3.34 5.74±3.58 0.014±0.05 0.020±0.03
Ours(w/o aux) 1.85±0.96 2.38±1.25 0.067±0.08 0.028±0.02
Ours (w/o align+aug+aux) 5.74±3.37 6.72±3.28 0.036±0.07 0.083±0.06
Ours 1.64±0.61 2.15±1.03 0.068±0.09 0.035±0.03
VII Conclusion and Feature Work
In this paper, we investigate the problem of robotic manipulation under limited observation and present a task-driven asynchronous active vision-action model. Our model first infers the optimal observation viewpoint accroding to the task goal and then determines the corresponding manipulation action based on the updated viewpoint’s obsevation. To enhance performance and stability, our model integrates three strategies: voxel alignment for spatial consistency, demonstration augmentation for better learning, and auxiliary rewards for real-time feedback. Together, these strategies help the model handle occlusions and improve sensor-motor coordination. Experimental results show that, compared to passive observation-based model, our model significantly improves manipulation performance through actively observation. It further learns sensor-motor coordination through interaction with the environment, rather than being restricted to a limited set of demonstration trajectories. Despite these promising results, the model’s performance can still be affected by complex occlusions and dynamic environments. Future research could explore incorporating real-time visual feedback to better handle dynamic changes in the environment.