### v2

# Observe Then Act: Asynchronous Active Vision-Action Model for Robotic Manipulation

Abstract
In real-world scenarios, many robotic manipulation tasks are hindered by occlusions and limited fields of view, posing significant challenges for passive observation-based models that rely on fixed or wrist-mounted cameras. In this paper, we investigate the problem of robotic manipulation under limited visual observation and propose a task-driven asynchronous active vision-action model. Our model serially connects a camera Next-Best-View(NBV) policy with a gripper Next-Best-Pose(NBP) policy, and trains them in a sensor-motor coordination framework using few-shot reinforcement learning. This approach enables the agent to reposition a third-person camera to actively observe the environment based on the task goal, and subsequently determine the appropriate manipulation actions. We trained and evaluated our model on 8 viewpoint-constrained tasks in RLBench. The results demonstrate that our model consistently outperforms baseline algorithms, showcasing its effectiveness in handling visual constraints in manipulation tasks. Our code and demonstration videos are available at https://hymwgk.github.io/ota/.
I INTRODUCTION
Although reinforcement learning-based models for robotic manipulation can learn to infer actions directly from image observations through interaction with the environment, they still heavily rely on carefully positioned static third-person or wrist-mounted cameras[1, 2, 3]. While these cameras reduce sample variance during interaction[3, 4], they often struggle with providing sufficient visual data in occluded or limited-visibility scenarios due to unadjustable viewpoints or fixed camera-gripper configuration.
However, humans can adjust their head and eye positions based on task goals. This helps them obtain advantageous visual information for action decision [5]. Inspired by this, recent active vision-action models enable agents to actively adjust their sensory viewpoints while interacting with the environment[3, 6, 7, 8], mimicking human capabilities.
Active vision-action models, like human perception, can be categorized as synchronous or asynchronous[10]. Synchronous models make decisions about visual and motor actions simultaneously based on current observations[7, 8, 11, 12]. However, the need for simultaneous coordination of sensor-motor actions places high demands on the model’s dynamic modeling capabilities and may limits its ability to handle tasks requiring large viewpoint shifts.
In contrast, asynchronous models decouple vision sensing and motor actions over time, allowing the agent to prioritize observation before interaction and then focus on inferring actions based on the selected viewpoint, thereby distributing sensor-motor coordination across the entire task episode. Although such models are commonly used for object grasping in occluded scenes[13, 14], we believe that the asynchronous approach may also be suitable for sequential manipulation tasks.
In this paper, we introduce a task-driven asynchronous active vision-action model for robotic manipulation under limited observation(Fig.LABEL:fig:head_image). Our model employs a sequential dual-agent structure, with a Next-Best-View(NBV) agent to infer optimal viewpoints and a Next-Best-Pose(NBP) agent to determine gripper actions based on the observations from the NBV-inferred viewpoints(Fig.3). By alternating between sensor and motor actions, the model actively perceives and interacts with the environment to achieve task objectives.
We train our model with few-shot reinforcement learning[2] and introduce a viewpoint-aware demo augmentation method to maximize the utilization of limited demonstration trajectories. To address the expanded observation-action space caused by the active viewpoint, we employ a viewpoint-centric voxel alignment technique that rotates the scene grid to align with the camera viewpoint. We also integrate a task-agnostic auxiliary reward to stabilize training, encouraging agents to interact effectively and select informative viewpoints for gripper action decisions. We evaluate our model on 8 RLBench[9] tasks with limited viewpoints, demonstrating its effectiveness in viewpoint selection and action inference based on task requirements. The main contributions of this paper are as follows:
  • We propose a task-driven asynchronous active vision-action model that uses a novel dual-agent system to achieve effective robotic manipulation under limited observation.
  • We introduce methodological innovations, including viewpoint-aware demo augmentation, viewpoint-centric voxel alignment, and a task-agnostic auxiliary reward strategy, to enhance model performance, learning efficiency, and training stability.
  • Comprehensive experiments on public RLBench tasks validate the effectiveness and robustness of our approach.
II Related Work
Active Vision for Manipulation. Active vision aims to enable agents to actively observe the environment based on specific task goals. It has been widely applied in tasks such as indoor visual navigation[12, 15], object searching[16, 17] and 3D reconstruction[18]. In robotic manipulation, active vision is primarily used for grasping pose detection in occluded environments[7, 13, 14, 19] and for detecting fruit nodes in automated harvesting[20, 21, 22]. Recent studies have extended the use of active vision to more general robotic manipulation tasks. For instance, Cheng et al[11] proposed an active vision model with a target detector module to handle object-pushing task in occluded scenes. Lv et al[8] introduced a model-based sensing-aware framework for manipulating flexible objects. Matthew et al[4] employed a self-prediction approach to learn viewpoint adjustments while simultaneously developing motion skills for robotic arm joints. Similarly, Shang et al[6] utilized a self-supervised reward mechanism to jointly learn visual sensing and motor skills. Unlike these methods, our active vision-aciton model does not require a task-specific target detector and employs an RL-based dual agent asynchronous decision-making structure.
3D Voxel-based Manipulation. 3D voxels are highly effective in representing spatial structures[23], making them widely used in domains such as medical imaging[24, 25] and autonomous driving [26, 27]. Recently, they have also been adopted in robotic manipulation models for visual representation. For instance, James et al.[2] transformed original RGB-D images into 11-dimensional voxel grids and developed a 3DUnet-based reinforcement learning model to extract scene features and infer manipulation actions. Similarly, Shridhar et al.[28] used discrete 3D voxel images as input, employing the Perceiver network[29] as a backbone for end-to-end behavioral cloning, trained on demonstration trajectories. Leveraging the strong spatial structure representation and robust structural invariance of voxels under viewpoint changes—which aids in viewpoint learning—our model also uses voxelized scene images as input.
Figure 2: Decision process comparison.We divide single step interaction into NBV for viewpoint and NBP for action selection. Reward $r_{\text{t'}}$ is derived from $r_{\text{t+1}}$ after NBP interaction, enabling sensor-motor joint training through shared task rewards.
Attention-driven Robotic Manipulation. The attention mechanism, widely applied in fields such as object detection and natural language processing, contributes to robotic manipulation by focusing on the most task-relevant scene information from raw image inputs, improving both training efficiency and model performance[30]. ARM [1] and C2FARM [2] use attention mechanisms to focus on key visual input regions, with ARM selecting noteworthy 2D regions via DQN, while C2FARM predicts regions of interest(ROIs) in 3D space. Liang et al. [31] pre-train a scene saliency attention encoder offline using human-annotated saliency maps and then train the reinforcement learning operation agent online based on the scene features output by this encoder. However, these methods largely rely on single or multiple fixed cameras from different viewpoints to capture images of the scene, lacking the capability to actively choose their viewpoints. Our model partially incorporates the attention mechanism from C2FARM and endows the model with the ability to predict the optimal task-related observation viewpoint, further mimicking the human’s active visual perception capability.
Figure 3: Overview of our model pipeline. The NBV policy infers the scene ROI position $\mathbf{f}^{*}_{t}$ based on the current observation and determines the optimal viewpoint $\mathbf{v}^{*}$ for observing the ROI to achieve the task goal. The NBP agent infers the gripper actions based on the scene ROI image from viewpoint $\mathbf{v}^{*}$ and the proprioceptive state.
III Problem Formulation
We divide one interaction time step $t\sim t+1$ into the viewpoint decision and transform process of the NBV policy $\pi_{\text{v}}$ from $t\sim t^{\prime}$ and the gripper action decision and interaction process of the NBP policy $\pi_{\text{p}}$ from $t^{\prime}\sim t+1$. The asynchronous active vision manipulation problem can be formalized as a POMDP described by the tuple $<\mathcal{O},\mathcal{A^{\text{c}}},P,\mathcal{A^{\text{g}}},P^{\prime},r,\gamma>$, as shown in Fig.2. where $\mathcal{O}$ is the observation space, $\mathcal{A^{\text{c}}},\mathcal{A^{\text{g}}}$ is the NBV and NBP action space, respectively. At time $t$, the NBV policy $\pi_{\text{v}}$ selects a camera action $\mathbf{a}_{t}^{\text{c}}=\pi_{\text{v}}(\cdot|\mathbf{o}_{t})\in\mathcal{A^{\text{c}}}$ based on the observation $\mathbf{o}_{t}\in\mathcal{O}$, resulting in a new observation $\mathbf{o}_{t^{\prime}}\in\mathcal{O^{\prime}}$ via transition probability $p(\mathbf{o}_{t^{\prime}}|\mathbf{o}_{t},\mathbf{a}_{t}^{\text{c}})\in P$. The NBP policy $\pi_{\text{p}}$ then selects a gripper action $\mathbf{a}_{t^{\prime}}^{\text{g}}=\pi_{\text{p}}(\cdot|\mathbf{o}_{t^{\prime}})\in\mathcal{A^{\text{g}}}$, with the next scene observation $\mathbf{o}_{t^{\prime}}$ obtained through $p^{\prime}(\mathbf{o}_{t+1}|\mathbf{o}_{t^{\prime}},\mathbf{a}_{t^{\prime}}^{\text{g}})\in P^{\prime}$. The environment provides a reward $r_{t+1}\in\mathbb{R}$. Our goal is to jointly learn $\pi^{*}_{\text{v}}$ and $\pi^{*}_{\text{p}}$ to maximize the expected sensor-motor reward $\mathbb{E}_{\pi_{\text{v}},\pi_{\text{p}}}\Big{[}\sum_{t=0}^{\infty}\gamma^{t}r(\mathbf{o}_{t},\mathbf{a}^{\text{c}}_{t},\mathbf{a}^{\text{g}}_{t^{\prime}})\Big{]}$.
Observation and Action Space. We define the observation spaces for the NBV and NBP policies as identical, with $\mathcal{O}=\{\mathcal{O}^{\text{v}},\mathcal{O}^{\text{p}}\}$, where $\mathbf{o}^{\text{v}}\in\mathcal{O}^{\text{v}}$ represents the 3D voxel grid derived from the RGB-D image, and $\mathbf{o}^{\text{p}}\in\mathcal{O}^{\text{p}}$ is the proprioceptive state. Specifically, the RGB-D image $\mathbf{o}^{\text{i}}$ is converted into a point cloud $\mathbf{o}^{\text{pc}}\in\mathbb{R}^{n\times 6}$ using the function $\mathbf{P}(\mathbf{o}^{\text{i}}):\mathbf{o}^{\text{i}}\rightarrow\mathbf{o}^{\text{pc}}$, and voxelized via $\mathbf{V}(\mathbf{o}^{\text{pc}},\mathbf{e},\mathbf{c}):\mathbf{o}^{\text{pc}}\rightarrow\mathbf{o}^{\text{v}}$, where $\mathbf{e}$ denotes the voxel resolution and $\mathbf{c}$ represents the voxel center. The resulting 3D voxel grid $\mathbf{o}^{\text{v}}\in\mathbb{R}^{n\times(3+M+1)}$ contains the center coordinates, $M$-dimensional features, and an occupancy flag for each voxel. The proprioceptive state $\mathbf{o}^{\text{p}}=\{\mathbf{v},\mathbf{g}\}$ comprises the camera viewpoint position $\mathbf{v}\in\mathbb{R}^{3}$ and the gripper pose $\mathbf{g}\in\mathbb{R}^{6}$.
Following common practice[14, 19], the camera’s movement is limited to the upper hemisphere at a fixed radius $r^{\text{v}}=1.2\text{m}$ above the table, with the optical axis directed at the sphere center, illustrated in Fig.8(b). The viewpoint is described by spherical coordinates $\mathbf{v}=(r^{\text{v}},\theta^{\text{v}},\phi^{\text{v}})$. The NBV policy action is $\mathbf{a}^{\text{c}}=\{\mathbf{k}^{\text{v}},\mathbf{k}^{\text{f}}\}$, where $\mathbf{k}^{\text{v}}\in\mathbb{R}^{2}$ represents the discrete viewpoint index for $(\theta^{\text{v}},\phi^{\text{v}})$, and $\mathbf{k}^{\text{f}}\in\mathbb{R}^{3}$ denotes the discrete center position index of a region of interest(ROI) that requires attention within the scene. The NBP policy action $\mathbf{a}^{\text{g}}=\{\mathbf{k}^{\text{g}},d\}$ consists of the discrete 6-DoF gripper pose index $\mathbf{k}^{\text{g}}$ and the goal distance $d$ between the gripper tips.
IV Method
IV-A Model Structure
We illustrates the basic pipline of our model in Fig3. Once the alignment process is completed(Section IV-B), the observation $\mathbf{o}_{t}$ is obtained. The DQN-based NBV policy network, parameterized by $\Theta$, then uses $\mathbf{o}_{t}$ to determine the discrete position index $\mathbf{k}^{\text{f}}$ of the ROI center in the scene, as described in Eq.(1). Simultaneously, the network infers the corresponding discrete index $\mathbf{k}^{\text{v}}$, which represents the viewpoint most likely to enhance task success. The camera is subsequently adjusted to the goal viewpoint $\mathbf{v}^{*}$, derived from the selected $\mathbf{k}^{\text{v}}$.
$\mathbf{a}_{t}^{\text{c}}=\arg\underset{\mathbf{k}^{\text{v}},\mathbf{k}^{\text{f}}}{\max}Q_{\Theta}(\mathbf{o}_{t},\mathbf{k}^{\text{v}},\mathbf{k}^{\text{f}})$ (1)
Based on the ROI position $\textbf{f}^{*}$, derived from $\mathbf{k}^{\text{f}}$, the image from $\mathbf{v}^{*}$ is cropped and aligned to generate the 3D voxel grid $\mathbf{o}^{\text{v}}_{t^{\prime}}$ for the ROI. The NBP policy network, also DQN-based and parameterized by $\Phi$, infers the gripper closure $d$ and pose index $\mathbf{k}^{\text{g}}$ based on $\mathbf{o}_{t^{\prime}}=\{\mathbf{o}^{\text{v}}_{t^{\prime}},\mathbf{o}^{\text{p}}_{t^{\prime}}\}$. Finally, a complete sensor-motor interaction step is achieved after gripper interaction and rewarded by $r_{t+1}$.
$\mathbf{a}_{t^{\prime}}^{\text{g}}=\arg\underset{\mathbf{k}^{\text{g}},d}{\max}Q_{\Phi}(\mathbf{o}_{t^{\prime}},\mathbf{k}^{\text{g}},d)$ (2)
We denote a complete interaction transition as $\text{T}_{t}=\{\text{T}_{t}^{\text{v}},\text{T}_{t}^{\text{p}},r_{t+1}\}$, where $\text{T}_{t}^{\text{v}}=(\mathbf{o}_{t},\mathbf{a}_{t}^{\text{c}},\mathbf{o}_{t^{\prime}})$ represents the NBV policy’s sub-transition, and $\text{T}_{t}^{\text{p}}=(\mathbf{o}_{t^{\prime}},\mathbf{a}_{t^{\prime}}^{\text{g}},\mathbf{o}_{t+1})$ represents the NBP policy’s sub-transition. After each training episode, transitions are stored in the replay buffer along with demonstration data. Both policies are updated using the same task reward $r_{t+1}$, thereby achieving sensor-motor joint learning. Each policy employs a separate 3D-CNN based U-Net[32] as the image encoder, followed by a three-layer fully connected network(FCN) for action decoding.
[…]
IV-B Viewpoint-centric Voxel Alignment
[…]
$\mathbf{o}^{\text{v}}=\mathbf{V}\big{(}(^{\text{L}}{\text{T}}_{\text{W}}\times\mathbf{o}^{\text{pc}}),\mathbf{e},\mathbf{c}\big{)}$ (3)
[…]
$\mathbf{v}^{*},\mathbf{f}^{*},\mathbf{g}^{*}=^{\text{W}}{\text{T}}_{\text{L}}\times(\mathbf{v}^{*}_{\text{L}},\mathbf{f}^{*}_{\text{L}},\mathbf{g}^{*}_{\text{L}})$ (4)
IV-C Viewpoint-aware Demo Augmentation
To provide early guidance and reduce exploration, we manually alternate camera and gripper movements to generate a small set of demonstration trajectories. Each demo trajectory $\tau_{\text{d}}=\{(\mathbf{o}_{i},\mathbf{a}_{i})\}^{N}$ consists of observation-action pairs collected at a fixed rate. To further leverage limited demo data, existing augmentation methods [1, 2, 28] first select key frame indices from the demo trajectory, construct transition samples from the key observation-action pairs, and then resample between key frames to generate additional transitions. However, they cannot handle the observation variance caused by viewpoint shifts. We define the key frame discovery function as $KF:\tau_{\text{d}}\rightarrow\big{\{}(k_{j}^{\text{c}},k_{j}^{\text{g}})\big{\}}^{M}\subset\mathbb{R}_{+}^{2}$, identifying key frames for both viewpoint and gripper movements. Gripper key frames $k_{j}^{\text{g}}$ are marked when the gripper is stationary or its closure state changes, while key viewpoints are defined as the first stable viewpoint between two gripper key frames, indexed by $k_{j}^{\text{c}}$. The key frame indices satisfy $10\leq k_{j-1}^{\text{c}}\leq k_{j-1}^{\text{g}}\leq k_{j}^{\text{c}}\leq k_{j}^{\text{g}}\leq N-$ 1. As shown in Fig.5, raw transition samples $\text{T}_{j}$ are then constructed as formalized in Eq.(5).
$\begin{split}\text{T}_{j}&=\{\text{T}_{j}^{\text{v}},\text{T}_{j}^{\text{p}},r_{j}\}\\ &=\{(\mathbf{o}_{k_{j-1}^{\text{g}}},\mathbf{a}_{k_{j-1}^{\text{g}}}^{\text{c}},\mathbf{o}_{k_{j}^{\text{c}}}),(\mathbf{o}_{k_{j}^{\text{c}}},\mathbf{a}_{k_{j}^{\text{c}}}^{\text{g}},\mathbf{o}_{k_{j}^{\text{g}}}),r_{j}\}\end{split}$ (5)
To generate additional augmented transitions, we uniformly sample within the intervals $k_{j-1}^{\text{g}}\sim k_{j}^{\text{c}}$ and $k_{j}^{\text{c}}\sim k_{j}^{\text{g}}$, obtaining the augmented indices$(k_{j-1}^{\text{g}})^{\text{aug}}$ and $(k_{j}^{\text{c}})^{\text{aug}}$. The resulting augmented transition $\text{T}_{j}^{\text{aug}}$ is formalized in Eq.(6).
$\begin{split}\text{T}_{j}^{\text{aug}}&=\{(\text{T}_{j}^{\text{v}})^{\text{aug}},(\text{T}_{j}^{\text{p}})^{\text{aug}},r_{j}\}\\ \end{split}$ (6)
Both $\text{T}_{j}$ and $\text{T}_{j}^{\text{aug}}$ are ultimately stored in the replay buffer as early training experiences for the model.
Figure 5: Demo trajectory augmentation.For each demo trajectory, we uniformly sample observations(represented by hatched circles) between both viewpoint key frames(solid blue circles) and gripper key frames(solid green circles) to generate additional transition samples, enabling viewpoint-aware demonstration augmentation.
IV-D Task-agnostic Auxiliary Reward
We use sparse reward to avoid designing task-specific rewards. However, since sparse reward only provides feedback at the end of an episode, it hinders efficient sensor-motor learning. Here we introduce two task-agnostic auxiliary rewards to provide real-time guidance to the training process.
To encourage interaction with the environment and accurate environment modeling, we propose the Interaction Encouragement Reward, denoted as $r_{\text{i}}$ As illustrated in Fig.6, $r_{\text{i}}$ provides positive signals when the NBV policy selects a reachable, non-empty ROI and penalizes actions targeting empty or unreachable areas due to execution failures. The specific value assignment for $r_{\text{i}}$ is detailed in Eq.(7).
$r_{\text{i}}=\begin{cases}0.02,&\text{ROI reachable and non-empty}\\ 0.0,&\text{ROI reachable but empty}\\ -0.02,&\text{ROI unreachable}\end{cases}$ (7)
To encourage the model to select viewpoints that capture more structural information from the ROI, we introduce the Entropy Reduction Reward($r_{\text{e}}$), inspired by [21]. The goal is to reduce occupancy entropy within the ROI after viewpoint adjustment, as shown in Fig.6. Assuming the ROI contains $N$ voxels, the occupancy entropy of the $i$-th voxel($e_{i}$) is calculated using the Shannon entropy function(Eq.(8)), where $p_{\text{o}}(i)$ represents its occupancy probability. If a voxel is occluded, $p_{\text{o}}$ is set to 0.5.
$e_{i}=-p_{\text{o}}(i)\log_{2}\big{(}p_{\text{o}}(i)\big{)}-\big{(}1-p_{\text{o}}(i)\big{)}\log_{2}\big{(}1-p_{\text{o}}(i)\big{)}$ (8)
The total occupancy entropy change of the ROI from time $t$ to $t^{\prime}$ is expressed as:
$\Delta E=\frac{1}{N}\sum_{i=1}^{N}e^{t}_{i}-\frac{1}{N}\sum_{i=1}^{N}e^{t^{\prime}}_{i}=E_{t}-E_{t^{\prime}}$ (9)
We define reward $r_{\text{e}}=k\cdot\Delta E$, where $k$ is empirically set to 1.0. The auxiliary reward $r_{\text{i}}$ requires cooperation between the NBV and NBP policies, while $r_{\text{e}}$ is exclusive to the NBV policy. Therefore, the reward for updating the NBP policy is expressed as $r_{\text{nbp}}=r_{\text{task}}+r_{\text{i}}$, and the NBV reward is $r_{\text{nbv}}=r_{\text{task}}+r_{\text{e}}+r_{\text{i}}$, where $r_{\text{task}}\in[0,1]$ is determined by task completion. These task-agnostic rewards generalize across tasks without the need for manual tuning.
Figure 6: Viewpoint adjustment and ROI selection in a 2D simplified schematic scene. The model is encouraged to select non-empty ROIs and reduce occluded voxels from the adjusted viewpoint $\mathbf{v}_{t^{\prime}}$.
V Experiment Setup
Task Setting. We trained and evaluated our model on RLBench[9], a widely-used public benchmark for robotic manipulation in reinforcement learning. Task selection followed two criteria: 1) each task must include at least one viewpoint with significant visual occlusion, and 2) no single fixed or wrist-mounted camera provides consistently low occlusion across selected tasks. As the ROI shifts throughout the manipulation process, and there may be phases without distinct objects(e.g., midair gripper movements), defining occlusion rates based on object segmentation, as done in previous methods[14, 21], is impractical. To address this, we define Task Occlusion Rate(TOR) for a given viewpoint as the average occupancy entropy of the ROI where the gripper’s goal actions occur, based on observations from that viewpoint across $L=100$ demo trajectories, as described in Eq.(10). Here, $E^{l}_{t^{\prime}_{k}}$ represents the ROI occupancy entropy at the $k$-th keyframe in trajectory $l$ after viewpoint adjustment where the ROI center is set to the gripper’s goal position in the demo observation. To compare the occlusion levels across selected tasks, we assume that the human-demoed active viewpoint contains the strongest task-specific prior knowledge. Thus, we can obtain the Relative TOR by dividing the TOR of the selected viewpoint by the TOR of the human-demoed active viewpoint(oracle).
$\text{TOR}=\frac{1}{L}\sum_{l=1}^{L}\sum_{k=0}^{K_{l}}E^{l}_{t_{k}^{\prime}}$ (10)
Fig.7 presents the relative TOR statistics for 8 selected RLbench tasks under overhead, front, wrist-mounted cameras as well as the human-demoed active camera.
Figure 7: Reletive task occlusion rate of selected tasks. We use the active viewpoint occlusion rate of 100 demonstration trajectories for each task as a benchmark.
Baselines. We compared our model against two baseline categories: 1) visual manipulation model that utilize static third-person and wrist-mounted cameras, and 2) active third-person camera-based models. For the first category, we compared our model against C2FARM [2] with static front, overhead and wrist-mounted camera settings to explore the potential effectiveness of our active vision-action model in handling tasks with limited observability. For the second category, we compared our model with Behavior Cloning(BC) to test whether it can learn new sensor-motor skills beyond demonstration imitation. We also tested a Random viewpoint model [3], which shares the same architecture as ours but with the NBV policy network disabled(learning rate set to zero). This setup allows us to assess how much performance improvement is achieved through active viewpoint adjustment as opposed to random choices.
Training Setting. We collected 30 demo trajectories for each task and used them as prior knowledge for training. C2FARM was implemented based on its official release and trained with sparse task rewards. The Behavior Cloning model was trained with MSE loss, and the Random model used the same reward structure as our model(SectionIV-D). All models were trained using Adam with a learning rate of $0.0005$ for 40,000 updates, with a batch size of 128. Experiments were conducted on a machine with 4 NVIDIA RTX 4090 GPUs.
TABLE I: Task success $\uparrow$ . Mean and std of 100 tests are reported(3 seeds).
Models Basketball in Hoop Meat off Grill Meat on Grill Open Drawer Press Switch Put Rubbish in Bin Take Money out Safe Unplug Charger On Average
C2FARM(wrist) 0.93±0.02 0.01±0.01 0.11±0.02 0.44±0.07 0.90±0.02 0.00±0.00 0.52±0.01 0.30±0.02 0.40±0.35
C2FARM(overhead) 0.87±0.02 0.00±0.00 0.28±0.05 0.03±0.01 0.89±0.03 0.61±0.04 0.24±0.03 0.89±0.04 0.48±0.36
C2FARM(front) 0.26±0.03 0.28±0.04 0.68±0.04 0.37±0.04 0.86±0.06 0.91±0.02 0.66±0.06 0.00±0.00 0.50±0.30
Random(active) 0.00±0.00 0.00±0.00 0.00±0.00 0.00±0.00 0.00±0.00 0.29±0.02 0.00±0.00 0.00±0.00 0.04±0.09
BC(active) 0.00±0.00 0.00±0.00 0.02±0.01 0.19±0.03 0.51±0.02 0.00±0.00 0.38±0.03 0.02±0.01 0.14±0.19
Ours(active) 0.81±0.02 0.56±0.04 0.73±0.04 0.50±0.03 0.85±0.03 0.96±0.01 0.78±0.02 0.92±0.01 0.76±0.15
TABLE II: Task episode $\downarrow$ . The maximum length of each task episode is limited to 10 steps.
Models Basketball in Hoop Meat off Grill Meat on Grill Open Drawer Press Switch Put Rubbish in Bin Take Money out Safe Unplug Charger On Average
C2FARM(wrist) 3.73±0.18 9.96±0.04 9.27±0.20 7.06±0.49 3.64±0.23 9.98±0.03 6.39±0.11 7.75±0.17 7.22±2.39
C2FARM(overhead) 5.21±0.17 10.00±0.01 8.50±0.28 9.78±0.07 4.13±0.23 6.52±0.21 8.46±0.21 3.93±0.22 7.07±2.30
C2FARM(front) 8.43±0.15 8.66±0.18 5.91±0.24 8.04±0.20 4.49±0.40 3.22±0.13 5.63±0.41 9.99±0.02 6.80±2.20
Random(active) 10.00±0.00 10.00±0.00 10.00±0.00 10.00±0.00 10.00±0.00 7.76±0.27 10.00±0.00 10.00±0.00 9.72±0.75
BC(active) 10.00±0.00 10.00±0.00 9.93±0.05 8.68±0.24 6.74±0.28 10.00±0.00 7.64±0.27 9.90±0.07 9.11±1.22
Ours(active) 5.35±0.08 6.43±0.28 5.15±0.26 6.65±0.19 3.89±0.28 3.33±0.09 4.30±0.17 3.91±0.07 4.88±1.16
VI Result
Main Performance. We first analyzed the average task success rate(SR) and episode length(EL) across different models to evaluate overall sensor-motor performance. Table I shows that our model achieved the highest success rate, outperforming the second-best by 21%, especially in occluded tasks like meat_on_grill, meat_off_grill and unplug_charger. Similarly, Table II shows that our model had the shortest average manipulation steps, outperforming the second-best by 28.2% and showing the least variance in step counts.
Fixed and wrist-mounted cameras performed well in low-occlusion tasks, such as basketball_in_hoop, press_switch, put_rubbish_in_bin, C2FARM occasionally matching or surpassing our model. However, their performance dropped sharply in occluded scenarios, leading to a higher variance. In contrast, our model remained robust in occluded tasks. Although the BC model demonstrated some ability to adjust its viewpoint(as detailed below), it struggled with limited training samples, achieving only a 14% success rate. The random viewpoint model performed the worst across all tasks, highlighting its inability to learn sensor-motor coordination.
Figure 8: Partial rollout cases and viewpoint distribution. (a) The image observations corresponding to the first three inferred viewpoints $\mathbf{v}^{*}$ and the gripper actions $\mathbf{g}^{*}$ for four representative tasks during the test episodes. (b) Red circles indicate manually set viewpoints from 30 demo trajectories, while blue triangles represent the learned viewpoint positions from successful episodes across 100 test runs.
Active Perception Strategy Analysis. We further analyzed the perception strategies of all models using two metrics: Average Single Step Information Gain(ASSIG) and Average Initial Information Gain(AIIG). ASSIG captures the average reduction in ROI entropy before and after a single viewpoint shift, while AIIG measures the entropy reduction compared to the initial viewpoint.
$\text{ASSIG}=\frac{1}{L}\sum_{l=1}^{L}\sum_{k=0}^{K_{l}}(E^{l}_{t_{k}}-E^{l}_{t^{\prime}_{k}})$ (11)
$\text{AIIG}=\frac{1}{L}\sum_{l=1}^{L}\sum_{k=0}^{K_{l}}([E^{l}_{t_{k}}]_{\mathbf{v}_{t_{0}}}-E^{l}_{t^{\prime}_{k}})$ (12)
Where $L=100$ denotes the number of testing episodes, $K_{l}$ represents the length of episode $l$, and $E^{l}_{t_{k}}$ is defined in Eq.(9). The term $[E^{l}_{t_{k}}]_{\mathbf{v}{t_{0}}}$ represents the visual entropy of the current ROI from the episode’s initial viewpoint. As shown in Table III, our model significantly enhances ROI observation by adjusting viewpoints, outperforming the BC model by 2.8% in ASSIG and 1.2% in AIIG. In contrast, static front and overhead cameras, which cannot shift viewpoints, fail to achieve information gain. The wrist-mounted camera, unable to move independently of the gripper, shows zero ASSIG and is unsuitable for AIIG computation.
Fig.8(b) compares the distribution of viewpoints learned by our model with the manually set active viewpoints from the demo trajectories used during training. Unlike demo patterns, our model’s learned viewpoints are more concentrated in some tasks(Basketball_in_hoop, unplug_charger) and more varied in others(meat_on_grill, open_drawer), indicating it learned to select task-relevant viewpoints rather than simply imitating demo data. The camera observations at key moments(Fig. 8) further showcase our model’s active visual perception capabilities.
Figure 9: Proportion of success, timeouts, and execution failures across test episodes for different models.
TABLE III: Comparison of information gain and interaction capacity across different models.
Models ASSIG $\uparrow$ (avg/max) AIIG $\uparrow$ (avg/max) First $\downarrow$ Interaction Non- $\downarrow$ interaction
C2FARM(wrist) 0.0/0.0 N/A 3.21±1.82 4.86±2.70
C2FARM(overhead) 0.0/0.0 0.0/0.0 2.97±1.70 4.41±2.77
C2FARM(front) 0.0/0.0 0.0/0.0 3.41±2.28 4.17±2.74
Random(active) 0.0/5.1% 0.0/5.2% 9.07±1.47 9.49±1.07
BC(active) 0.7%/2.3% 5.6%/18.1% 6.31±2.46 7.54±2.15
Ours(active) 3.5%/6.9% 6.8%/25.1% 1.64±0.61 2.15±1.03
Environmental Modeling and Interactive Capabilities. We evaluated the models’ environment modeling and interaction capabilities by recording success, timeout(TO), and execution failure(EF) rates over 100 test episodes(Fig.9). The C2FARM model exhibited a high failure rate(up to 42%) due to trajectory execution failures. Similarly, the random viewpoint model saw the majority of its test episodes end in failure because of execution errors. Although only 4% of the BC model’s episodes ended due to execution errors, 82% failed due to timeouts. In contrast, our model showed a 13% failure rate from trajectory errors and only 9% from timeouts, indicating effective integration of low-level motion execution.
As shown in Table III, we also analyzed the average steps to first interaction(FI) and non-interactive steps(NI) per model. The BC model exhibited frequent non-interactive steps, correlating with its high timeout rate, suggesting difficulty learning precise manipulations from limited demonstrations. In contrast, our model initiated interactions more quickly, reducing non-interactive steps, and demonstrating superior accuracy in modeling environmental dynamics and structural understanding after training.
Ablation Study. We evaluated the impact of three strategies on manipulation performance:(1) viewpoint-centric voxel alignment(Section IV-B), (2) viewpoint-aware demo augmentation(Section IV-C), and (3) task-agnostic auxiliary rewards(Section IV-D). Separate ablation tests were conducted for each, as well as a combined test removing all strategies.
Table IV shows that each strategy effectively improved success rates(SR), with voxel alignment and viewpoint-aware augmentation contributing the most, as success rates dropped by approximately 40% without them and auxiliary rewards added around 6% improvement.
Beyond success rate improvements, all three strategies contributed to reductions in episode length(EL), timeouts rate(TO), execution failures(EF), and the average number of non-interactive steps(NI) per episode. This demonstrates their broader importance to the model’s effectiveness. Interestingly, while the voxel alignment strategy significantly boosts overall performance, we found it leads to a reduction in ASSIG score. We hypothesize that without viewpoint alignment, the NBV agent avoids potential issues with out-of-bounds actions and clipping, which may simplify the learning of viewpoint transformations. However, the overall performance still confirms that viewpoint alignment plays a crucial role in improving operational effectiveness.
Ablating all strategies led to a sharp decline, reducing the average success rate to 22%, underscoring the synergistic effect of combining these strategies for optimal performance.
TABLE IV: Ablation over voxel alignment,demo augmentation and task-agnostic auxiliary rewards.
Ablation Setting Result Proportion Interactivity Sensing
align aug aux SR EF TO EL FI NI AIIG ASSIG
0.34 0.53 0.14 7.61 4.33 5.81 0.046 0.082
0.31 0.55 0.15 7.99 4.49 5.74 0.014 0.020
0.70 0.16 0.15 5.52 1.85 2.38 0.067 0.028
0.22 0.57 0.22 8.46 5.74 6.72 0.036 0.083
0.76 0.15 0.09 4.88 1.64 2.15 0.068 0.035
VII Conclusion and Future Work
In this paper, we investigate the problem of robotic manipulation under limited observation and present a task-driven asynchronous active vision-action model. Our model first infers the optimal viewpoint based on the task goal, then selects the manipulation action from the updated observation. To enhance performance and training stability, we integrate voxel alignment for spatial consistency, demonstration augmentation for enhanced learning, and auxiliary rewards for real-time training guidance. These strategies enable the model to handle occlusions and improve sensor-motor coordination. Experimental results show that our model significantly outperforms passive observation-based models by actively refining viewpoints and learning sensor-motor coordination through interaction with the environment. However, complex occlusions and dynamic environments remain challenging. Future work could explore real-time visual feedback to better adapt to environmental changes.